OpenStudy (anonymous):

stuck on equation Write the equation of the line passing through the points (-6,1) and (-4,2).

OpenStudy (anonymous):

Write the equation of the line passing through the points (-6,1) and (-4,2). m = 2 - 1/ -4 - (-6) = 1/ 2 = 0.5 y - 1 = 0.5 [x - (-6) ] y -1 = 0.5 [ x + 6] y - 1 =

OpenStudy (anonymous):

Start with finding the slope of your line.

OpenStudy (anonymous):

Then plug into the point slope formula and simplify.

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (anonymous):

You're slope of 1/2 is correct.

OpenStudy (anonymous):

thats what i did im stuck i've done other ones like this

OpenStudy (anonymous):

just confused cuz the one is negative

OpenStudy (anonymous):

i just want to know if im doing it right

OpenStudy (anonymous):

\[y-1=0.5(x+6)\]

OpenStudy (anonymous):

is this correct ? m= 2 -(-4)/ -2 - 0 = 6/ -2 = -3 y - (-4) = -3 (x-0) y + 4 = -3x + 0 y + 4 - 4 = -3x + 0 - 4 y = -3x - 4

OpenStudy (anonymous):

wrong equation srry

OpenStudy (anonymous):

is this correct m = 2 - 1/ -4 - (-6) = 1/ 2 = 0.5 y - 1 = 0.5 [x - (-6) ] y -1 = 0.5 [ x + 6] y - 1 + 1 = x + 6 + 1 y = x + 7

OpenStudy (anonymous):

Close, but not correct.

OpenStudy (anonymous):

\[y−1=0.5(x+6)\]

OpenStudy (anonymous):

\[y-1= 0.5x+3\]

OpenStudy (anonymous):

\[y=0.5x+4\]

OpenStudy (anonymous):

where do you get the 0.5x+3

OpenStudy (anonymous):

I added the 1 over to the other side to give you y =

OpenStudy (anonymous):

no how did u get y−1=0.5(x+6) to y−1=0.5x+3

OpenStudy (anonymous):

Hold on.

OpenStudy (anonymous):

k

OpenStudy (anonymous):